OpenStudy (anonymous):

I WILL GIVE YOU A MEDAL. ' ' 1.A backpack has a mass of 50 kg. what is it's weight on earth? 2.A satellite with a mass of 400kg has a weight of 4000N on a newly discovered moon of saturn. What is the gravitational strength of the newly discovered moon? 3.Calculate the forece on a toy helicopter that has a mass of 12kg and an acceleration of 4m/s^2

OpenStudy (michele_laino):

1) Weigh W of an object, is given by the subsequent formula: W=m*g where m is the mass of that object, and g is the gravity, namely, g=9.81 m/sec^2

OpenStudy (michele_laino):

2) the requested gravitational strength G, is given by the subsequent ratio: \[G = \frac{F}{M}\] where F= 4000 Newtons, and M=400 Kg

OpenStudy (michele_laino):

3) According to the second principle of Newton, we have that the requested force F, is given by the subsequent formula: \[F = m \times a\] where: m=12 Kg, and a= 4 m/sec^2